# Consensus formation in quality-sensitive interdependent agent systems

### v2

@@ Abstract
We perform a comprehensive analysis of a collective decision-making model inspired by honeybee behavior. This model integrates individual exploration for option discovery and social interactions for information sharing, while also considering option qualities. Our assessment of the decision process outcome employs standard consensus metrics and investigates its correlation with convergence time, revealing common trade-offs between speed and accuracy. Furthermore, we investigate fluctuations arising from finite-size effects. In scenarios with strong interdependence, we identify an absorbing phase transition and scrutinize its key features, including critical fluctuations in finite spatial systems. We find that fluctuations play a significant role in the collective decision-making process, particularly when options have similar qualities. Our findings are especially relevant for finite adaptive systems, as they provide insights into navigating decision-making scenarios with similar options more effectively. Finally, we show that this model also fulfills Weber’s Law of perception, aligning with previous analyses of collective-decision problems in their adherence to different psychophysical laws.
@@ I Introduction
Collective decision making is an emergent, self-organized phenomenon by which a group of agents, each with its own decision mechanisms at the individual level, reaches an agreement on a certain task or topic [1, 2, 3, 4]. From humans taking part in elections, social mammal herds, flocking fish, foraging insect colonies or robot swarms, collective decision making is an ubiquitous process across all scales of biological and artificial complexity [1, 2, 5, 6, 7]. Thus, addressing this question requires interdisciplinary approaches integrating fields such as sociology, behavioral ecology, biology, physics, computer science and communication studies, among many others [1, 8]. Through empirical observations, field experiments, laboratory studies, and computational modeling, researchers strive to unravel the intricate dynamics of collective decision making and its significance.
Many efforts have been devoted in the modeling of opinion dynamics in order to understand how a collective makes a decision among a set of options. Initial attempts devised models that encompassed simple imitation rules controlling the build up of an agreement between equivalent options, such as the voter model [9, 10] or the majority rule model [11]. This models, though, have been regarded as too unrealistic in their basic formulation, and many more complex models have been proposed in order to capture a richer spectrum of social behaviors, such as partisanship, heterogeneity, personalized information or non linear interactions [12, 13, 14, 15, 16, 17].
@@
In recent years, the collective behavior exhibited by social insects has become a focal point of interest within the scientific community. They conceive a paradigmatic example of systems formed by rather simple units exhibiting this kind of collective emergent phenomena. In particular, the nest-site selection process in honeybee swarms, widely studied by biologists and ecologists [18, 19, 20], has laid the ground for a great number of collective decision making models inspired in this fascinating process [21, 22, 23, 24, 25], and for further exploration of these dynamics in robot swarms [26, 27, 28, 29].
@@
In a bee-like house-hunting process, agents can adopt a wider spectrum of behaviors beyond the usual imitation rules. Initially, they must discover available options either through individual exploration or by communicating with peers using different signaling mechanisms. Additionally, they have the ability to abandon options and become uncommitted or neutral. Lastly, they may cross-inhibit other peers with different opinions. Importantly, in honeybee-inspired models, agents are able to estimate the qualities of options, which significantly impacts their subsequent signaling behavior [30]. Specifically, the better they estimate the quality to be, the longer they will advertise this option to their peers. This mechanism allows a positive feedback loop to regulate the group’s opinion formation, discarding the worst option in favor of the more suitable ones. Ultimately, a collective agreement is reached through a combination of individual exploration of the environment to identify possible options, estimation of their quality, and information exchange among members of the group.
@@
List, Estholtz and Seeley presented an agent-based model (in the following, the LES model) of the decision making process performed by honeybee swarms [21], further developed analytically later by T. Galla [31], which focuses on these three features mentioned above. The LES model is a simple model that introduces option qualities into the decision process. By balancing individual option exploration and social interactions, it effectively captures decision-making dynamics. The main result highlighted in the original investigation is that when individuals possess sufficiently accurate estimates of the qualities and engage in a high degree of social interaction, the swarm tends to converge towards collective agreement, particularly when faced with various options of differing quality values.
@@
In this paper, we delve deeper into studying the formation of consensus by inspecting the LES model. Our investigation focuses on the interplay between social feedback and individual exploration, which the model balances through a single parameter known as the “interdependence” parameter. The intensity of individual, independent exploration serves a dual purpose. On one hand, it is essential for the group to perceive all available options or to adapt in changing environments [28, 32]. On the other hand, excessive individual exploration can introduce noise that hampers the effectiveness of the aforementioned positive feedback loop. Agents tend to overlook the actual state of their peers and adopt opinions based on the outcome of their own exploration. Consequently, stronger social interaction is required to balance this effect. Recent work highlights the capability of decentralized systems to effectively balance noise with social interactions under various scenarios [33, 34, 35].
@@
We approach the problem by considering the mean-field model equations and complementing these results with simulations in fully connected networks and regular lattices in finite dimensions. These simulations help us address finite size effects and convergence times. We conduct an exhaustive parametric exploration of the model, which intriguingly has led us to establish a connection between a specific limit of this model and the well-known contact process. Following recent developments in collective decision-making models inspired by honeybee behavior, we demonstrate how this model may also adhere to certain psychophysical laws that describe organisms’ perception when discriminating between various stimuli or options [36].
The paper is organized as follows. In Section II, we provide a brief review of the model formulation and its analytical solution, highlighting results in specific limiting scenarios. Section III presents a thorough analysis of the model and its significant findings. In Sections III.1 and III.2, we initiate our investigation by examining the possibility of consensus formation and its characteristic convergence time within a binary decision context. This is followed by an exploration of the significance of finite size effects in Section III.3. In Section III.4, we investigate the correlation between the equal-quality and equal-discovery binary model and the contact process, along with its characteristic non-equilibrium critical behavior. Furthermore, in Section III.5, we discuss the model’s alignment with Weber’s Psychophysical Law of perception. We close this section by analyzing a multi-site extension of the model that encompasses more options than the typical binary scenario (Section III.6). Finally in Section IV, we present our main conclusions.
@@ II Modeling fully-connected scout-bee-like swarms
List, Elsholtz, and Seeley introduced their honeybee-inspired collective decision-making model in [21]. This model encapsulates the primary mechanisms through which a swarm of honeybees achieves a collective decision. Honeybees possess the ability to independently discover potential nest site options and evaluate their quality. Subsequently, they communicate their findings to the rest of the swarm. Concurrently, bees can interact with their peers and opt to adopt an opinion already present within the group.
@@
The uncommitment transition probabilities ${r_{\alpha,t}}$ are inherently linked to the qualities of the sites. In the original LES model, transitions from committed to uncommitted states occur deterministically after a predetermined commitment time has elapsed. An agent commits to an option and advertises it for a fixed duration. To render the model’s equations mathematically tractable, T. Galla [31] replaced this deterministic process by a stochastic process defined by the following rates:
@@
The stochastic problem can be analyzed using a master equation, from which one can derive a set of nonlinear differential equations describing the evolution of the average fraction values $\langle f_{\alpha,t}\rangle$, henceforth referred to as the average dancing frequencies, for each state $\alpha$. By following the mathematical details outlined in [31], and assuming a fully connected, mean-field-like system, one can arrive at the following equations:
@@ II.1.3 Limit of equal-quality sites
In the specific case where we have equal-quality sites and fully interdependent bees ($\pi_{\alpha}=0$ for all $\alpha$), we find two solutions $f_{0}^{*}=1$ and $f_{0}^{*}=r/\lambda$ delimited by the stability threshold value $\lambda=r$. In the former solution, all the other frequencies vanish, i.e. $f_{\alpha}^{*}=0$ for $1\leq\alpha\leq k$, whereas in the latter the actual values of $f_{\alpha}^{*}$ remain undetermined (note that a null denominator appears in Eq. (5)). According to LSA, any combination that satisfies $\sum_{\alpha=1}^{k}f_{\alpha}=1-r/\lambda$ is a potential solution. Consequently, the stationary state reached after numerical integration of Eq. (4) or in numerical simulations of the model with those parameters is highly dependent on the initial conditions and is susceptible to significant finite-size effects (see Secs. III.3 and III.4).
@@ III Results
Going beyond the mean-field approach, previous studies have delved into the stochastic dynamics of this decision-making model through simulations conducted on a regular lattice [31] and on random networks [29]. Additionally, experiments have been conducted using mini-robots as a physical platform [29], albeit with a limited parameter set explored. In the following, we focus on characterizing the model’s behavior using a mean-field approach, conducting an exhaustive exploration across a broad parameter range. To quantify the stationary results, we rely on analytical expressions. Simulations conducted on fully connected systems or square lattices yield average results that align perfectly with these analytical solutions. However, to complement our analysis, we take advantage of simulations to measure the time needed to reach the stationary state and to address finite-size effects.
@@
We will first concentrate on the specific case of a binary decision problem, a topic extensively studied in various general opinion dynamics models [11, 12, 10, 37, 38, 14, 39, 40, 41, 17]. This scenario has also been the focus of other honeybee-inspired models [22, 23, 24, 25]. The generalization to larger values of $k$ is discussed later in Section III.6.
@@
Figure 4: A,B: Color-maps illustrating the consensus parameters $Q_{2}$ and $Q_{2}^{\prime}$, as defined in Eqs. (11) and (12) respectively with $x=2$, across the parameter space $(\pi_{1,2},\lambda$). The black lines represent the crossover points $Q_{2}=0$ or $Q_{2}^{\prime}=0$. C: Time to reach the stationary state, $t_{ss}$, obtained from simulations with a swarm size of $N=5000$ bees. The fully interdependent limit ($\pi_{\alpha}=0$ or $\lambda=1$) is not included in the maps (see discussion in Sec.III.4). The site quality parameters are held constant at $(q_{1},q_{2})=(7,10)$ across all three maps. D,E: Crossover lines $\lambda^{*}(Q_{2},Q_{2}^{\prime}=0)$ as a function of $\pi_{1,2}$. Each curve represents a different value of $q_{1}$, while $q_{2}=10$ is held constant. F: Time to reach the stationary state, $t_{ss}$, obtained from simulations plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$, for three values of the interdependence. Parameters $\pi_{1}=\pi_{2}=0.1$ and $N=5000$ are fixed. Simulations start from uncommitted initial conditions. The markers are color-coded based on the consensus established at the stationary state. They are slightly jittered around the $q_{1}$ axis to allow for better visualization of the color code.
@@
As detailed in the model description section, we consistently assume that site 2 holds the highest quality, denoted by $q_{2}>q_{1}$. Additionally, we assume that bees independently assess the qualities of the sites, represented by the model parameter $\mu=0$. This implies that the abandonment rates of the dances depend solely on the qualities of each site, defined as $r_{\alpha}=q_{0}/q_{\alpha}$. Previous research on this model has shown that, under these conditions, the swarm can identify the best site across a wide range of interdependence values, $\lambda$ [21, 31]. In this setup, we investigate the interplay between group communication, represented by the model parameter $\lambda$, and individual exploration success, captured by $\pi_{\alpha}$. We explore scenarios where sites may vary in their likelihood of being discovered, denoted by $\pi_{1}\neq\pi_{2}$ (asymmetric scenario), or where they share the same discovery probability, denoted by $\pi_{1}=\pi_{2}$ (symmetric scenario).
@@ III.1 Symmetric discovery scenario
While setting a threshold value on one of the populations provides a straightforward measure of consensus, this method may fail to capture scenarios where the population committed to the best option is substantial, yet faces significant competition from other sites. Additionally, it might overlook situations where one site clearly leads, but the committed population has not yet met the prescribed threshold. To address these concerns, we adopt the consensus definition proposed by LES in their original paper  [21]. This approach involves comparing the two largest values of $f_{\alpha}$, offering a more comprehensive assessment of the decision-making dynamics. In a general manner, we can define a consensus measure as follows:
@@
We would also like to acknowledge that the definition of consensus in Eq. (11) may not be very appropriate to describe the low discovery region of the parameter space where $\pi_{1,2}\rightarrow 0$. There, we observe that $\lambda^{*}$ decreases, suggesting that a scarce environment with low $\pi_{1,2}$ is the most beneficial situation for the swarm, as it allows for the widest range of $\lambda$ yielding consensus. However, it’s important to note that the range of uncommitted population increases drastically in that region, and the system can even transition to an absorbing state around $\lambda=r_{2}$. Therefore, this effect should be taken into account in the quorum measure. In the original paper by LES [21], the consensus definition is strengthened by requiring a minimum proportion of agents engaged in the decision process. Here, we propose a slightly different definition to encompass this effect along with the site competition.
@@
The trend observed is similar to that of consensus: interdependence tends to increase the decision time across all values of $\pi_{1,2}$, while the discovery probabilities speed up the process, reducing $t_{ss}$. This correlation between the established consensus value and the time to reach the stationary state suggests a speed-accuracy trade-off, commonly observed in biological systems [42] an in models of collective decision making [27, 41]. Achieving the best decision typically takes longer at higher values of $\lambda$, while processes are quicker below the consensus crossover line, where a less accurate consensus, such as a simple majority, can be reached much faster. Similarly, increasing the value of the discovery probabilities, and thus the introduction of independent information, accelerates the decision process at the cost of a lower value of $Q^{\prime}$, which can be even negative, depending on $\lambda$. Only at the limits $\lambda\rightarrow 0$ or $\lambda\rightarrow 1$ $\pi_{1,2}$ has practically no effect: in the former, there’s no discussion mediated by interdependence, resulting in a quickly established yet inaccurate stationary state, while in the latter, social feedback dominates the process, with little impact from individual exploration. It’s interesting to note the agreement between the stationary times discussed here and the relaxation times obtained from the linear stability analysis eigenvalues, as depicted in Fig. A1.

### v3

# Consensus formation and relative stimulus perception in quality-sensitive, interdependent agent systems

@@ Abstract
We perform a comprehensive analysis of a collective decision-making model inspired by honeybee behavior. This model integrates individual exploration for option discovery and social interactions for information sharing, while also considering option qualities. Our assessment of the decision process outcome employs standard consensus metrics and investigates its correlation with convergence time, revealing common trade-offs between speed and accuracy. Furthermore, we show the model’s compliance with Weber’s Law of relative stimulus perception, aligning with previous analysis of collective decision problems. Our study also identifies non-equilibrium critical behavior in specific limits of the model, where the highest values of consensus are achieved. This result highlights the intriguing relationship between optimal performance, critically, and the fluctuations caused by finite size effects, often seen in biological systems. Our findings are especially relevant for finite adaptive systems, as they provide insights into navigating decision-making scenarios with similar options more effectively.
@@ I Introduction
Collective decision making is an emergent, self-organized phenomenon by which a group of agents, each with its own decision mechanisms at the individual level, reaches an agreement on a certain task or topic Dyer et al. (2009); Krause and Ruxton (2002); Baronchelli (2018); Bose et al. (2017). From humans taking part in elections, social mammal herds, flocking fish, foraging insect colonies or robot swarms, collective decision making is an ubiquitous process across all scales of biological and artificial complexity Dyer et al. (2009); Krause and Ruxton (2002); Sumpter (2010); Smith and Harper (2003); Leonard et al. (2024). Thus, addressing this question requires interdisciplinary approaches integrating fields such as sociology, behavioral ecology, biology, physics, computer science and communication studies, among many others Dyer et al. (2009); Sasaki and Pratt (2018). Through empirical observations, field experiments, laboratory studies, and computational modeling, researchers strive to unravel the intricate dynamics of collective decision making and its significance.
Many efforts have been devoted in the modeling of opinion dynamics in order to understand how a collective makes a decision among a set of options. Initial attempts devised models that encompassed simple imitation rules controlling the build up of an agreement between equivalent options, such as the voter model Holley and Liggett (1975); Castellano et al. (2009a) or the majority rule model Galam, S. (2002). This models, though, have been regarded as too unrealistic in their basic formulation, and many more complex models have been proposed in order to capture a richer spectrum of social behaviors, such as partisanship, heterogeneity, personalized information or non linear interactions Galam (2008); Castellano et al. (2009b); Redner (2019); De Marzo et al. (2020); Bizyaeva et al. (2023); Ramirez et al. (2024).
@@
In recent years, the collective behavior exhibited by social insects has become a focal point of interest within the scientific community. They conceive a paradigmatic example of systems formed by rather simple units exhibiting this kind of collective emergent phenomena. In particular, the nest-site selection process in honeybee swarms, widely studied by biologists and ecologists Britton et al. (2002); Seeley and Buhrman (2001); Seeley et al. (2012), has laid the ground for a great number of collective decision making models inspired in this fascinating process List et al. (2008); Valentini et al. (2014); Reina et al. (2015, 2017); Gray et al. (2018), and for further exploration of these dynamics in robot swarms Reina et al. (2018a); Valentini et al. (2016); Talamali et al. (2021); March-Pons et al. (2024a).
@@
In a bee-like house-hunting process, agents can adopt a wider spectrum of behaviors beyond the usual imitation rules. Initially, they must discover available options either through individual exploration or by communicating with peers using different signaling mechanisms. Additionally, they have the ability to abandon options and become uncommitted or neutral. Lastly, they may cross-inhibit other peers with different opinions. Importantly, in honeybee-inspired models, agents are able to estimate the qualities of options, which significantly impacts their subsequent signaling behavior Passino et al. (2008). Specifically, the better they estimate the quality to be, the longer they will advertise this option to their peers. This mechanism allows a positive feedback loop to regulate the group’s opinion formation, discarding the worst option in favor of the more suitable ones. Ultimately, a collective agreement is reached through a combination of individual exploration of the environment to identify possible options, estimation of their quality, and information exchange among members of the group.
@@
List, Estholtz and Seeley presented an agent-based model (in the following, the LES model) of the decision making process performed by honeybee swarms List et al. (2008), further developed analytically later by T. Galla Galla (2010), which focuses on these three features mentioned above. The LES model is a simple model that introduces option qualities into the decision process. By balancing individual option exploration and social interactions, it effectively captures decision-making dynamics. The main result highlighted in the original investigation is that when individuals possess sufficiently accurate estimates of the qualities and engage in a high degree of social interaction, the swarm tends to converge towards collective agreement, particularly when faced with various options of differing quality values.
@@
In this paper, we delve deeper into studying the formation of consensus by inspecting the LES model. Our investigation focuses on the interplay between social feedback and individual exploration, which the model balances through a single parameter known as the “interdependence” parameter. The intensity of individual, independent exploration serves a dual purpose. On one hand, it is essential for the group to perceive all available options or to adapt in changing environments Talamali et al. (2021); Aust et al. (2022). On the other hand, excessive individual exploration can introduce noise that hampers the effectiveness of the aforementioned positive feedback loop. Agents tend to overlook the actual state of their peers and adopt opinions based on the outcome of their own exploration. Consequently, stronger social interaction is required to balance this effect. Recent work highlights the capability of decentralized systems to effectively balance noise with social interactions under various scenarios Zakir et al. (2022); Khaluf et al. (2018); Rausch et al. (2019).
@@
Building on recent advances in collective decision-making models inspired by honeybee behavior, we show how the LES model adheres to Weber’s law of perception Reina et al. (2018b); Pais et al. (2013). In the context of decision-making, Weber’s law states that the smallest detectable difference between two options is proportional to the magnitude of their average quality. This means that as stimulus intensity increases, a greater difference is needed to perceive a change. The law, which applies to various sensory modalities such as vision, hearing, and touch, illustrates the logarithmic nature of perception, where sensitivity diminishes as stimulus intensity rises. Remarkably, our findings reveal that varying levels of social feedback, exploration, and system size can influence the minimum noticeable quality difference in collective decision-making.
Finally, we demonstrate that a limiting case of the LES model maps directly onto the well-known contact process. Our results near this limit reveal how the bio-inspired system reaches maximum consensus at the onset of critical behavior, illustrating how such systems may exploit criticality to enhance consensus formation - similar to how many biological systems are hypothesized to operate near a critical point Mora and Bialek (2011); Muñoz (2018); Chialvo et al. (2020).
We approach the study of the LES model by analyzing its mean-field equations and complementing these results with simulations on fully connected networks and regular lattices in finite dimensions. These simulations allow us to explore the impact of finite size effects and examine convergence times, providing a more comprehensive understanding of the model’s behavior in both idealized and realistic settings.
The paper is organized as follows. In Section II, we provide a brief review of the model formulation and its analytical solution, highlighting results in specific limiting scenarios. Section III presents a thorough analysis of the model and its significant findings. In Sections III.1 and III.2, we initiate our investigation by examining the possibility of consensus formation and its characteristic convergence time within a binary decision context. In Section III.3, we show the model’s adherence to Weber’s Psychophysical Law of perception. Following this, in Section III.4, we explore the significance of finite size effects, which leads us to identify, in Section III.5, a correlation between the equal-quality and equal-discovery binary model and the contact process, highlighting its characteristic non-equilibrium critical behavior. Finally, we extend the model to incorporate more than two options in a multi-site scenario in Section III.6, before summarizing and concluding our study in Section IV.
@@ II Modeling fully-connected scout-bee-like swarms
List, Elsholtz, and Seeley introduced their honeybee-inspired collective decision-making model in List et al. (2008). This model encapsulates the primary mechanisms through which a swarm of honeybees achieves a collective decision. Honeybees possess the ability to independently discover potential nest site options and evaluate their quality. Subsequently, they communicate their findings to the rest of the swarm. Concurrently, bees can interact with their peers and opt to adopt an opinion already present within the group.
@@
The uncommitment transition probabilities ${r_{\alpha,t}}$ are inherently linked to the qualities of the sites. In the original LES model, transitions from committed to uncommitted states occur deterministically after a predetermined commitment time has elapsed. An agent commits to an option and advertises it for a fixed duration. To render the model’s equations mathematically tractable, T. Galla Galla (2010) replaced this deterministic process by a stochastic process defined by the following rates:
@@
The stochastic problem can be analyzed using a master equation, from which one can derive a set of nonlinear differential equations describing the evolution of the average fraction values $\langle f_{\alpha,t}\rangle$, henceforth referred to as the average dancing frequencies, for each state $\alpha$. By following the mathematical details outlined in Galla (2010), and assuming a fully connected, mean-field-like system, one can arrive at the following equations:
@@ II.1.3 Limit of equal-quality sites
In the specific case where we have equal-quality sites and fully interdependent bees ($\pi_{\alpha}=0$ for all $\alpha$), we find two solutions $f_{0}^{*}=1$ and $f_{0}^{*}=r/\lambda$ delimited by the stability threshold value $\lambda=r$. In the former solution, all the other frequencies vanish, i.e. $f_{\alpha}^{*}=0$ for $1\leq\alpha\leq k$, whereas in the latter the actual values of $f_{\alpha}^{*}$ remain undetermined (note that a null denominator appears in Eq. (5)). According to LSA, any combination that satisfies $\sum_{\alpha=1}^{k}f_{\alpha}=1-r/\lambda$ is a potential solution. Consequently, the stationary state reached after numerically integrating Eq. (4) is highly sensitive to the initial conditions. Any initial state satisfying $\sum_{\alpha=1}^{k}f_{\alpha}=1-r/\lambda$ will remain unchanged, while other initial conditions that do not satisfy this condition will evolve toward stationary points where one option dominates, and the others vanish. The specific option that prevails depends on which one initially has an advantage. When performing numerical simulations, finite-size effects further influence the decision outcome (see Secs. III.4 and III.5).
@@ III Results
Going beyond the mean-field approach, previous studies have explored the stochastic dynamics of this decision-making model through simulations on a regular lattice Galla (2010) and on random networks March-Pons et al. (2024a). Additionally, experiments have been conducted using mini-robots as a physical platform March-Pons et al. (2024a), albeit with a limited parameter set explored. In these random scenarios, the deviation from mean-field behavior is primarily influenced by network connectivity. As analyzed in March-Pons et al. (2024a), the emergence of a giant component within the communication network is crucial for the system to transition toward mean-field-like results.
@@
In the following, we focus on characterizing the model’s behavior using a mean-field approach, conducting an exhaustive exploration across a broad parameter range. To quantify the stationary results, we rely on analytical expressions. Simulations conducted on fully connected systems or square lattices yield average results that align perfectly with these analytical solutions. However, to complement our analysis, we take advantage of simulations to measure the time needed to reach the stationary state and to address finite-size effects.
@@
We will first concentrate on the specific case of a binary decision problem, a topic extensively studied in various general opinion dynamics models Galam, S. (2002); Galam (2008); Castellano et al. (2009a); Starnini et al. (2012); Valentini et al. (2017); Redner (2019); Vazquez et al. (2019); Ramirez et al. (2022); Reina et al. (2024); Ramirez et al. (2024). This scenario has also been the focus of other honeybee-inspired models Valentini et al. (2014); Reina et al. (2015, 2017); Gray et al. (2018). The generalization to larger values of $k$ is discussed later in Section III.6.
@@
Figure 4: A,B: Color-maps illustrating the consensus parameters $Q_{2}$ and $Q_{2}^{\prime}$, as defined in Eqs. (11) and (12) respectively with $x=2$, across the parameter space $(\pi_{1,2},\lambda$). The black lines represent the crossover points $Q_{2}=0$ or $Q_{2}^{\prime}=0$. C: Time to reach the stationary state, $t_{ss}$, obtained from simulations with a swarm size of $N=5000$ bees. The fully interdependent limit ($\pi_{\alpha}=0$ or $\lambda=1$) is not included in the maps (see discussion in Sec.III.5). The site quality parameters are held constant at $(q_{1},q_{2})=(7,10)$ across all three maps. D,E: Crossover lines $\lambda^{*}(Q_{2},Q_{2}^{\prime}=0)$ as a function of $\pi_{1,2}$. Each curve represents a different value of $q_{1}$, while $q_{2}=10$ is held constant. F: Time to reach the stationary state, $t_{ss}$, obtained from simulations plotted against the quality of the inferior option, $q_{1}$, while maintaining $q_{2}=10$, for three values of the interdependence. Parameters $\pi_{1}=\pi_{2}=0.1$ and $N=5000$ are fixed. Simulations start from uncommitted initial conditions. The markers are color-coded based on the consensus established at the stationary state. They are slightly jittered around the $q_{1}$ axis to allow for better visualization of the color code.
@@
As detailed in the model description section, we consistently assume that site 2 holds the highest quality, denoted by $q_{2}>q_{1}$. Additionally, we assume that bees independently assess the qualities of the sites, represented by the model parameter $\mu=0$. This implies that the abandonment rates of the dances depend solely on the qualities of each site, defined as $r_{\alpha}=q_{0}/q_{\alpha}$. Previous research on this model has shown that, under these conditions, the swarm can identify the best site across a wide range of interdependence values, $\lambda$ List et al. (2008); Galla (2010). In this setup, we investigate the interplay between group communication, represented by the model parameter $\lambda$, and individual exploration success, captured by $\pi_{\alpha}$. We explore scenarios where sites may vary in their likelihood of being discovered, denoted by $\pi_{1}\neq\pi_{2}$ (asymmetric scenario), or where they share the same discovery probability, denoted by $\pi_{1}=\pi_{2}$ (symmetric scenario).
@@ III.1 Symmetric discovery scenario
While setting a threshold value on one of the populations provides a straightforward measure of consensus, this method may fail to capture scenarios where the population committed to the best option is substantial, yet faces significant competition from other sites. Additionally, it might overlook situations where one site clearly leads, but the committed population has not yet met the prescribed threshold. To address these concerns, we adopt the consensus definition proposed by LES in their original paper  List et al. (2008). This approach involves comparing the two largest values of $f_{\alpha}$, offering a more comprehensive assessment of the decision-making dynamics. In a general manner, we can define a consensus measure as follows:
@@
We would also like to acknowledge that the definition of consensus in Eq. (11) may not be very appropriate to describe the low discovery region of the parameter space where $\pi_{1,2}\rightarrow 0$. There, we observe that $\lambda^{*}$ decreases, suggesting that a scarce environment with low $\pi_{1,2}$ is the most beneficial situation for the swarm, as it allows for the widest range of $\lambda$ yielding consensus. However, it’s important to note that the range of uncommitted population increases drastically in that region, and the system can even transition to an absorbing state around $\lambda=r_{2}$. Therefore, this effect should be taken into account in the quorum measure. In the original paper by LES List et al. (2008), the consensus definition is strengthened by requiring a minimum proportion of agents engaged in the decision process. Here, we propose a slightly different definition to encompass this effect along with the site competition.
@@
The trend observed is similar to that of consensus: interdependence tends to increase the decision time across all values of $\pi_{1,2}$, while the discovery probabilities speed up the process, reducing $t_{ss}$. This correlation between the established consensus value and the time to reach the stationary state suggests a speed-accuracy trade-off, commonly observed in biological systems Chittka et al. (2009) an in models of collective decision making Valentini et al. (2016); Reina et al. (2024). Achieving the best decision typically takes longer at higher values of $\lambda$, while processes are quicker below the consensus crossover line, where a less accurate consensus, such as a simple majority, can be reached much faster. Similarly, increasing the value of the discovery probabilities, and thus the introduction of independent information, accelerates the decision process at the cost of a lower value of $Q^{\prime}$, which can be even negative, depending on $\lambda$. Only at the limits $\lambda\rightarrow 0$ or $\lambda\rightarrow 1$ $\pi_{1,2}$ has practically no effect: in the former, there’s no discussion mediated by interdependence, resulting in a quickly established yet inaccurate stationary state, while in the latter, social feedback dominates the process, with little impact from individual exploration. It’s interesting to note the agreement between the stationary times discussed here and the relaxation times obtained from the linear stability analysis eigenvalues, as depicted in Fig. A1.